It takes 2 boys 2.5 hours to build a fort. How long would it take 3 boys to build 2 such forts?
Answer by ptaylor(2198)

It takes 2 boys 2.5 hours to build a fort. How long would it take 3 boys to build 2 such forts?
Let x=amount of time it takes all three boys working together to build 2 forts
Each fort takes 2*2.5 or 5 boy-hours to build. This means that one boy can build a fort in 5 hours.
In other words, each boys works at the rate of 1/5 forts per hour
Together, the three boys work at the rate of 1/5 + 1/5 +1/5=3/5 forts per hour.
So our equation to solve is:
(3/5)*x=2 (two forts, that is) multiply each side by 5
3x=10 divide each side by 3
x=3 1/3 hours---amount of time it takes all three working together
Another way to look at this problem:
If it takes 5 boy-hours to build 1 fort, then it takes 10 boy-hours to build 2 forts. With three boys working, it would take 10/3=3 1/3 hours
CK
(3/5)(10/3)=6/3=2 forts
Hope this helps---ptaylor